OpenStudy (anonymous):

Make: x the subject of D = n / x^3

jhonyy9 (jhonyy9):

Dx*3 =n x*3=n/D x=+/-(n/D)*(1/3)

OpenStudy (anonymous):

Thank you !

jhonyy9 (jhonyy9):

my pleasure - push the good answer please